the ph of an aqueous solution of Ba(OH)2 is 10 if the ksp of Ba(OH)2 is 1*10^-9 then the concentration of Ba+2 ions in the aqueous solution is

Grade:12

1 Answers

Roshan Mohanty
11 years ago

[OH(-)] = 10^(-4)
2.[OH(-)] = 2*10^(-4)

Ba(OH)2 <=> Ba(2+) + 2.OH(-)
Ksp=10^-9

so,
[Ba(2+)].(2.10^-4)^2 = 10^-9
[Ba(2+)] = 0.25 * 10^-1 (Ans)

Please Approve. . .
Thank You Smile